# PreAlign: Boosting Cross-Lingual Transfer by Early Establishment of Multilingual Alignment

### v1

@@ 5 Experiments on Synthetic Transferring Settings
5.1 Investigating Cross-Lingual Transfer based on Cloned English
We construct a synthetic language called En-Clone, by cloning all English words by a one-to-one mapping. En-Clone shares the same linguistic properties with English, such as vocabulary distribution, grammar and syntax, yet they have no word overlapping. See Figure 3 for an illustration of the creation of En-Clone.
@@ 5.1 Investigating Cross-Lingual Transfer based on Cloned English
5.2 Experimental Results
We present the results on LM, ZS-CLT and CLKA in Table 1. Beside Joint Training and PreAlign, we also list the performance of Only-Tgt, where we only train the model on the same amount of En-clone data, and Full-Tgt, where we train the model on the En-clone data with the same size as full English data.
Table 1 shows that compared to Only Tgt, Joint training achieves notable improvements on LM despite there are neither parallel signal or pivot words between English and English-clone. However, this transfer does not work well on CLKA, which is consistent with previous findings (Gao et al., 2024) that CLKA cannot be improved by multilingual pretraining.
PreAlign improves over Joint Training on all evaluation tasks.
We can also see that PreAlign significantly outperforms Joint Training on all three evaluation tasks. On the LM evaluation, PreAlign even achieves performance comparable to Full Tgt, despite it only uses 1% En-Clone data. This demonstrates the effectiveness of PreAlign for facilitating cross-lingual transfer.
5.3 An in-depth investigation of CLKA
In order to better investigate the dynamic of cross-lingual knowledge transfer during pretraining, we plot the accuracy of knowledge completion of different training period. Figure 4 presents the results.
Language ability affects the rate of knowledge learning.
Firstly, we can see from the top-left of Figure 4, where we test English knowledge in English language, models’ knowledge completion accuracy after each learning period rapidly grows as the pretraining goes on. This indicates that the rate of knowledge learning strongly correlates with models’ language modeling ability. The final performance also correlates with the knowledge frequency in the learning period as expected.
In the top-right of Figure 4 where we test English-Clone knowledge in English-clone language, we observe a similar trend with the top-left figure. However, the growing rate of Joint Training is slower compared to PreAlign especially when the frequency of knowledge is low, indicating the early alignment introduced by PreAlign can boost target language modeling ability, therefore improving the learning of target language knowledge.
PreAlign unlocks cross-lingual knowledge transfer.
From the bottom two figures in Figure 4, we can see the CLKA ability of Joint Training is significantly weaker than PreAlign. This renders PreAlign a promising method for learning truly multilingual knowledge alignment.
5.4 Ablation Study
In this section, we present an ablation study of the proposed methods, including the multilingual alignment initialization and the input-only codeswitching strategy. The results is presented in Table 2.
@@ 5.4 Ablation Study
LM Codeswitching Ratio
@@
Table 3: Comparison of the original codeswitching strategy and the proposed input-only codeswitching strategy. Not the codeswitching ratio in the table refers to the portion of random English samples that contains English-clone scripts during inferencing.
@@
          $\beta=25\%$ 17.0 78.2 58.5
          $\beta=50\%$ 16.8 78.6 60.9
          $\beta=75\%$ 16.6 78.8 62.1
          $\beta=100\%$ 16.5 79.3 64.6
@@
5.5 Multilingual alignment is maintained across pretraining.
@@
Joint Training 20.3 79.8 32.5 34.8 39.6 82.3 65.8 65.3 56.9 63.7 - 31.2 30.5 34.1 29.7
PreAlign 19.9 75.2 28.3 30.7 33.6 82.4 70.0 69.3 65.6 68.2 - 50.2 51.0 49.3 48.9
@@
Joint Training 15.8 62.2 24.0 27.7 31.2 84.3 70.8 70.6 63.7 68.6 - 36.7 35.6 36.4 33.0
PreAlign 16.1 58.0 23.3 25.3 29.4 83.9 74.0 72.9 68.2 71.4 - 54.3 53.1 52.4 50.1
Table 5: Performance on LM, ZS-CLT and CLKA of Joint Training and PreAlign across different scale of models.
In order to understand how the injected multilingual alignment information before pretraining evolves, we compute the similarity of aligned word embeddings at different training steps. Figure 5 illustrates the results.
Firstly, we can see that despite there are no vocabulary overlap between English and English-clone, the embedding similarity of aligned words still grows during pretraining, which is consistent with findings in previous works Dufter and Schütze (2020). This indicates the ability of spontaneous establishment of multilingual alignment of language models. Secondly, the aligned similarity score of PreAlign is near perfect as designed, and despite the score decreases at the beginning of pretraining, it maintains to be significantly higher than Joint Training throughout the pretraining process. Finally, the codeswitching strategy is helpful for both Joint Training and PreAlign, as it accelerates the increment of Joint Training’s aligning similarity score, and helps slow down the decrement of PreAlign’s aligning similarity score.
@@ 6 Experiments on Real-world Settings
We have presented experiments on a synthetic English to English-Clone settings. In this section, we aim to validate the effectiveness of PreAlign under real-world settings. Specifically, we consider the transfer from English to Chinese, Russian, German and Arabia. The target languages spans four different language families and serves as good representatives of world languages. Performances of LM, ZS-CLT and CLKA is shown in Table 5.
@@
It can be seen from Table 5 that PreAlign can still achieve substantially better performance compared to the original Joint Training method. This improvements is consistent across different model scales, rendering the effectiveness of PreAlign in real-world scenarios.
@@ Acknowledgement
Shujian Huang are the corresponding authors. This work is supported by National Science Foundation of China (No. 62376116, 62176120).

### v2

@@ 5.2 In-Depth Investigation for CLKA
Knowledge learning ability correlates with language ability.
We can see from the top-left of Figure 4, where we test English knowledge in English language, models’ knowledge completion accuracy after each learning period rapidly grows as the pretraining goes on. This indicates that the models’ ability to acquire knowledge correlates with their language modeling ability. The final performance also correlates with the knowledge frequency in the learning period as expected.
In the top-right of Figure 4 where we test English-Clone knowledge in English-clone language, we observe a similar trend as the top-left figure. However, the growing rate of PreAlign is higher compared to Joint Training especially when frequency of knowledge is low, thanks to better transfer of language ability from English to English-Clone.
PreAlign unlocks CLKA.
From the bottom two figures in Figure 4, we can see the CLKA ability of Joint Training is greatly weaker than PreAlign, close to the random guessing performance. This renders PreAlign a promising method for learning truly multilingual knowledge alignment.
5.3 Ablation Study
@@
In this section, we present an ablation study of the proposed methods. The results are in Table 3.
@@
Input-only codeswitching causes less mixed-script problem.
LM Mixed-Script Ratio
@@
Table 4: Comparison of the original codeswitching strategy and the proposed input-only codeswitching strategy. Note the mixed-script ratio in the table refers to the portion of random English samples that contains English-clone scripts during inference.
@@
          $\beta=25\%$ 17.0 78.2 80.2
          $\beta=50\%$ 16.8 78.6 83.1
          $\beta=75\%$ 16.6 78.8 88.4
          $\beta=100\%$ 16.5 79.3 90.3
@@
5.4 Maintaining Multilingual Alignment across Pretraining.
@@
Joint Training 20.3 79.8 32.5 34.8 39.6 82.3 65.8 65.3 56.9 63.7 - 37.8 39.5 36.1 37.7
PreAlign 19.9 75.2 28.3 30.7 33.6 82.4 70.0 69.3 65.6 68.2 - 63.8 66.5 64.7 63.6
@@
Joint Training 15.8 62.2 24.0 27.7 31.2 84.3 70.8 70.6 63.7 68.6 - 49.6 44.1 45.5 48.0
PreAlign 16.1 58.0 23.3 25.3 29.4 83.9 74.0 72.9 68.2 71.4 - 71.1 73.9 72.7 72.5
Table 6: Performance of Joint Training and PreAlign across different scale of models on language modeling, zero-shot cross-lingual transfer (ZS-CLT) and cross-lingual knowledge application (CLKA).
In order to understand how the injected multilingual alignment information evolves during pretraining, we compute the similarity of aligned word embedding at different training period (every 250 training steps). Figure 5 illustrates the results.
Firstly, we can see that despite there are no vocabulary overlap between English and English-clone, the embedding similarity of aligned words still grows during Joint-Training, which is consistent with findings in previous works Dufter and Schütze (2020). This indicates the ability of spontaneous establishment of multilingual alignment of language models. Secondly, the aligned similarity score of PreAlign is near perfect as designed, and despite the score decreases at the beginning of pretraining, it maintains to be significantly higher than Joint-Training throughout the pretraining process. Finally, the codeswitching strategy is helpful for both Joint Training and PreAlign, as it accelerates the increment of Joint Training’s aligning similarity score, and helps slow down the decrement of PreAlign’s aligning similarity score.
@@ 6 Experiments on Real-world Settings
We validate the effectiveness of PreAlign under real-world settings. Performances of LM, ZS-CLT and CLKA is shown in Table 6.
@@
It can be seen from Table 6 that PreAlign can still achieve substantially better performance compared to the original Joint Training method. This improvements is consistent across different model scales, rendering the effectiveness of PreAlign in real-world scenarios. Interestingly, the transferring effect from English is more preeminent for German and Russian than Chinese and Arabic, indicating typological similarity between language might also play important roles in cross-lingual transferring effectiveness.
@@ Acknowledgement
We would like to thank the anonymous reviewers for their insightful comments. Shujian Huang is the corresponding author. This work is supported by National Science Foundation of China (No. 62376116, 62176120) and Nanjing University-China Mobile Communications Group Co.,Ltd. Joint Institute.
Appendix A Experimental Details
Collection of the multilingual table
We recognize words using the word_tokenize function from NLTK library. The word set consists of all words, including named entities, that appears above 20 times in the pretraining corpus. We translate all words using GPT-4, by asking it to generate 5 most common translations of a given word (without placing it in a context).
Training details
During the multilingual alignment stage, we set the $\tau$ to be $0.1$. During the language pretraining stage, we independently sample sentences for LM loss and word pairs for the alignment loss at each training step. We ran all experiments on 8×8$\times$8 ×A100 GPUs. The multilingual alignment stage takes about 500 steps, and the language pretraining stage takes about 24000 steps. The running time of different sizes of models ranges from 4 hours to 24 hours.
Example of the synthesized knowledge
We collect relations from WikiData, and ask GPT-4 to compose templates for each relation. We then fill in the person name to synthesize knowledge about people. For example, if the subject, relation and object are Oprah Winfrey, godparent and Tyler Perry respectively, then the composed knowledge is Oprah Winfrey is the godparent of Tyler Perry.
En De Zh
150M
Joint Training 90.6 71.5 78.9
PreAlign 90.1 76.1 83.7
400M
Joint Training 92.6 75.1 83.3
PreAlign 92.3 78.9 85.6
1.3B
Joint Training 94.1 79.7 85.2
PreAlign 94.3 82.4 87.9
Table 7: ZS-CLT performance of Joint Training and PreAlign across different scale of models on the PAWSX dataset.
Appendix B Experimental results on the PAWSX dataset
To further validate the effectiveness of PreAlign, we conduct additional experiments on the PAWSX dataset (Yang et al., 2019), and show the result in Table 7. It can be seen that PreAlign can still bring consistent improvements across different model scales.